# Non-stationary Gaussian random fields on hypersurfaces: Sampling and strong error analysis

### v1

@@ 3.3. Error analysis of the SFEM discretization
$E_{1}$ $=\|\big{(}\tilde{\gamma}(\mathcal{L})-\tilde{\gamma}(\mathcal{L}_{h})P_{h}\big{)}\mathcal{L}^{-\eta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
$E_{2}$ $=\|\tilde{\gamma}(\mathcal{L}_{h})\big{(}P_{h}\mathcal{L}^{-\eta/2}\mathcal{W}-\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}\big{)}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
For the term $E_{1}$, note that by Bonito et al. (2024, Lemma 4.2),
$\|\mathcal{L}^{-\eta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq\frac{\eta}{\eta-d/2}.$
In particular, $\mathcal{L}^{-\eta/2}\mathcal{W}$ is in $L^{2}(\Omega;L^{2}(\mathcal{M}))$ and an application of Proposition 2.3 yields that
$E_{1}\lesssim C_{\zeta}(h)h^{\min\{\zeta;2\}},$
where
$C_{\zeta}(h)=\begin{cases}1,&\zeta\neq 2,\\ |\log h|,&\zeta=2.\\ \end{cases}$
For $E_{2}$, we first write $P_{h}\mathcal{L}^{-\eta/2}\mathcal{W}=\sum_{i=1}^{N_{h}}\beta_{i}e_{i}^{h}$, where $\beta_{i}=(P_{h}\mathcal{L}^{-\eta/2}\mathcal{W},e_{i}^{h})_{L^{2}(\mathcal{M})}=(\mathcal{L}^{-\eta/2}\mathcal{W},e_{i}^{h})_{L^{2}(\mathcal{M})}=\sum_{k=1}^{\infty}\lambda_{k}^{-\eta/2}W_{k}(e_{k},e_{i}^{h})_{L^{2}(\mathcal{M})}$. Further, following the definition of $\tilde{\mathcal{W}}$ in Equation 34, we have $\mathcal{L}_{h}^{-\eta/2}P_{h}\mathcal{W}=\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{-\eta/2}\xi_{i}e_{i}^{h}$ where $\xi_{i}=(e_{i}^{h},\mathcal{W})_{L^{2}(\mathcal{M})}=\sum_{k=1}^{\infty}W_{k}(e_{i}^{h},e_{k})_{L^{2}(\mathcal{M})}$. Note in particular that both $\beta_{i}$ and $\xi_{i}$ are real-valued (since the eigenbases $\{e_{k}\}_{k\in\amsmathbb{N}}$ and $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ are real-valued). Thus,
@@
$E_{2}^{2}=\sum_{i=1}^{N_{h}}|\tilde{\gamma}(\lambda_{i}^{h})|^{2}\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}],$
where we use the orthonormality of the eigenfunctions. Note that
$\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}]=\operatorname{\amsmathbb{E}}[|\beta_{i}|^{2}]-2\operatorname{\amsmathbb{E}}[\beta_{i}(\lambda_{i}^{h})^{-\eta/2}{\xi_{i}}]+\operatorname{\amsmathbb{E}}[(\lambda_{i}^{h})^{-\eta}|\xi_{i}|^{2}].$
First, by the independence of the sequence $\{W_{k}\}_{k=1}^{\infty}$ and $(\lambda_{k},e_{k})$ eigenpairs of $\mathcal{L}$,
$\operatorname{\amsmathbb{E}}[|\beta_{i}|^{2}]$ $=\sum_{k=1}^{\infty}\lambda_{k}^{-\eta}|(e_{k},e_{i}^{h})|^{2}_{L^{2}(\mathcal{M})}=\left(\sum_{k=1}^{\infty}\lambda_{k}^{-\eta}(e_{i}^{h},e_{k})_{L^{2}(\mathcal{M})}e_{k},e_{i}^{h}\right)_{L^{2}(\mathcal{M})}$
$=(\mathcal{L}^{-\eta}e_{i}^{h},e_{i}^{h})_{L^{2}(\mathcal{M})}=\|\mathcal{L}^{-\eta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}.$
Similarly, and using the fact that $(\lambda_{i}^{h},e_{i}^{h})$ is an eigenpair of $\mathcal{L}_{h}$,
$\operatorname{\amsmathbb{E}}[(\lambda_{i}^{h})^{-\eta}|\xi_{i}|^{2}]=(\lambda_{i}^{h})^{-\eta}\sum_{k=1}^{\infty}|(e_{i}^{h},e_{k})|^{2}_{L^{2}(\mathcal{M})}=(\lambda_{i}^{h})^{-\eta}\|e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}=\|\mathcal{L}_{h}^{-\eta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}$
and
$\operatorname{\amsmathbb{E}}[\beta_{i}(\lambda_{i}^{h})^{-\eta/2}{\xi_{i}}]=(\lambda_{i}^{h})^{-\eta/2}\sum_{k=1}^{\infty}\lambda_{k}^{-\eta/2}|(e_{k},e_{i}^{h})|^{2}=(\mathcal{L}^{-\eta/2}e_{i}^{h},\mathcal{L}^{-\eta/2}_{h}e_{i}^{h})_{L^{2}(\mathcal{M})}.$
This implies that by Proposition 2.3,
$\operatorname{\amsmathbb{E}}[|\beta_{i}-(\lambda_{i}^{h})^{-\eta/2}\xi_{i}|^{2}]=\|\mathcal{L}^{-\eta/2}e_{i}^{h}-\mathcal{L}^{-\eta/2}_{h}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}\lesssim h^{\min\{2\eta;4\}}C_{\eta}(h)^{2},$
where
$C_{\eta}(h)=\begin{cases}1,&\eta\neq 2,\\ |\log h|,&\eta=2.\end{cases}$
Hence, it holds that
$E_{2}^{2}\lesssim h^{\min\{2\eta;4\}}C_{\eta}(h)^{2}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}.$
It remains to bound $\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}$. To this end, recall Equation 12 and Equation 6 which yield that
$\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\lesssim\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{-\zeta}\lesssim\sum_{i=1}^{N_{h}}i^{-2\zeta/d}\lesssim\begin{cases}N_{h}^{1-2\zeta/d},&2\zeta/d<1,\\ \log(N_{h}),&2\zeta/d=1,\\ 1,&2\zeta/d>1,\end{cases}\lesssim\begin{cases}h^{-2(d/2-\zeta)},&\zeta<d/2,\\ |\log(h)|,&\zeta=d/2,\\ 1,&\zeta>d/2,\end{cases}$
where we used in the last inequality that $N_{h}\propto h^{-d}$. We obtain
$\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\lesssim h^{2(\zeta-\max\{\zeta;d/2\})}K_{\zeta}(h)^{2}$
with
$K_{\zeta}(h)=\begin{cases}1,&\zeta\neq d/2\\ |\log(h)|^{1/2},&\zeta=d/2\end{cases}$
We conclude that
$E_{2}\lesssim C_{\eta}(h)K_{\zeta}(h)h^{\min\{\eta;2\}+\zeta-\max\{\zeta;d/2\}}$
and in particular,
$h^{-\min\{\zeta;2\}}E_{2}\lesssim C_{\eta}(h)K_{\zeta}(h)h^{\beta},$
where we take $\beta=\min\{\eta;2\}+\zeta-\max\{\zeta;d/2\}-\min\{\zeta;2\}$. We have two cases. On the one hand, if $\zeta\geq 2$ then, since $\eta>\zeta$, we also have $\eta>2$ and therefore $\beta=0$, $C_{\eta}(h)=1$ and $K_{\zeta}(h)=1$ (given that $d/2<2$). On the other hand, if $\zeta<2$, then $\max\{\zeta;d/2\}<2$ and $\max\{\zeta;d/2\}<\eta$ and therefore $\beta=\min\{\eta;2\}-\max\{\zeta;d/2\}>0$. In turn, this means that $C_{\eta}(h)K_{\zeta}(h)h^{\beta}\lesssim|\log(h)|^{3/2}|h^{\beta}\lesssim 1$. Hence, in both cases
$h^{-\min\{\zeta;2\}}E_{2}\lesssim 1,$
and we conclude that
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\lesssim h^{\min\{\zeta;2\}}\left(C_{\zeta}(h)+h^{-\min\{\zeta;2\}}E_{2}\right)\lesssim C_{\zeta}(h)h^{\min\{\zeta;2\}}.$
The result follows by inserting the definition of $\zeta$. ∎
@@
For the error between $\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}$ and $\widehat{\mathsf{Z}}_{h}$ we get the following estimate, inspired by Bonito et al. (2024, Lemma 4.4).
@@
Let $\mathcal{E}_{h}=\big{\|}\left(\tilde{\mathcal{Z}}_{h}\right)^{-\ell}-\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}=\big{\|}\left(\gamma(\mathcal{L}_{h})\widetilde{\mathcal{W}}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\widetilde{\mathcal{W}}^{-\ell})\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}$. We note that $\widetilde{\mathcal{W}}$ is an $S_{h}^{\ell}$-valued random variable. Therefore, we can apply Equation 24 to realizations of $\widetilde{\mathcal{W}}$, and take the expectation on both sides to get
@@
We then distinguish two cases. First, if $d=1$, then we note that $\min\{\alpha+d/4;1\}\geq\alpha+d/4>d/2$ since $\alpha>d/4$. Besides, Bonito et al. (2024, Lemma 4.2) yields that for all $r\in(d/2,2)$, there is a constant $C>0$ such that
@@
If now $d=2$, since $\alpha+d/4>d/2=1$, we have $\min\{\alpha+d/4;1\}=1$. We then note that by the proof of Bonito et al. (2024, Lemma 4.2), for any $\varepsilon>0$
@@
It holds that there is a constant $C>0$ such that
$\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\leq Ch^{2}.$
@@
where Dziuk and Elliott (2013, Lemma 4.1) was applied in the final step. We conclude that

### v2

@@ 3.3. Error analysis of the SFEM discretization
$S_{1}$ $=\|\gamma(\mathcal{L})\mathcal{W}-\gamma(\mathcal{L})P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},\quad S_{2}=\|\gamma(\mathcal{L})P_{h}\mathcal{W}-\gamma(\mathcal{L}_{h})P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
To bound the term $S_{1}$, we use
$S_{1}^{2}$ $=\|\gamma(\mathcal{L})(I-P_{h})\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=\|\sum_{i\in\amsmathbb{N}}\gamma(\lambda_{i})W_{i}(I-P_{h})e_{i}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
$=\sum_{i\in\amsmathbb{N}}\gamma(\lambda_{i})^{2}\|(I-P_{h})e_{i}\|_{L^{2}(\mathcal{M})}^{2},$
and since $\gamma$ is an $\alpha$-amplitude spectral density, we get
(41) $S_{1}^{2}\lesssim\sum_{i\in\amsmathbb{N}}\lambda_{i}^{-2\alpha}\|(I-P_{h})e_{i}\|_{L^{2}(\mathcal{M})}^{2},$
Applying the Bramble–Hilbert lemma (69) with $t=2((\alpha-d/4)-\varepsilon_{h})\in(0,2)$ together with Equation 41 gives
$S_{1}^{2}\lesssim\sum_{i\in\amsmathbb{N}}h^{2t}\lambda_{i}^{-2\alpha+t}=h^{4(\alpha-d/4-\varepsilon_{h})}\sum_{i\in\amsmathbb{N}}\lambda_{i}^{-2(d/4+\varepsilon_{h})}\lesssim h^{4(\alpha-d/4)}\sum_{i\in\amsmathbb{N}}\lambda_{i}^{-2(d/4+\varepsilon_{h})},$
since $h^{-4\varepsilon_{h}}=e^{4(\alpha-d/4)}\lesssim 1$. We can then write, using Equation 6,
$S_{1}^{2}$ $\lesssim h^{4(\alpha-d/4)}\sum_{i\in\amsmathbb{N}}i^{-(1+4\varepsilon_{h}/d)}\lesssim h^{4(\alpha-d/4)}\int_{1}^{\infty}t^{-(1+4\varepsilon_{h}/d)}\,\mathrm{d}t$
$\lesssim\frac{1}{4\varepsilon_{h}/d}h^{4(\alpha-d/4)}\lesssim|\log h|h^{4(\alpha-d/4)},$
where we used the definition of $\varepsilon_{h}$ for the last inequality. So, we conclude $S_{1}\lesssim|\log h|^{1/2}h^{2(\alpha-d/4)}$.
For the term $S_{2}$, we note that
$S_{2}=\|\gamma(\mathcal{L})P_{h}\mathcal{W}-\gamma(\mathcal{L}_{h})P_{h}(P_{h}\mathcal{W})\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
where, following Equation 38, $\|P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=\|\sum_{j=1}^{N_{h}}\xi_{j}e_{j}^{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}=N_{h}<\infty$ (cf. Remark 3.2). Hence, we can use Proposition 2.4 to deduce that since $\alpha\leq 1$,
$S_{2}\lesssim|\log h|h^{2\alpha}\|P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
Note then that since $N_{h}\propto h^{-d}$ (uniform mesh assumption), we can write $\|P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}=N_{h}^{1/2}\lesssim h^{-d/2}$. Therefore, we can conclude that
$S_{2}\lesssim|\log h|h^{2\alpha-d/2}.$
Putting together the bounds obtained for $S_{1}$ and $S_{2}$, we can then conclude that if $\alpha\leq 1$,
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $\leq S_{1}+S_{2}\lesssim|\log h|h^{2\alpha-d/2}=|\log h|h^{2\min\{(\alpha-d/4);1\}},$
where we use the fact that $\alpha-d/4<\alpha<1$ to derive the last inequality.
Case $\alpha>1$
We now assume that $\alpha>1$. We define the function $\tilde{\gamma}(x)=\gamma(x)x^{\zeta/2}$. Note that $\gamma$ is an $\alpha$-amplitude spectral density and $\tilde{\gamma}$ decays as $|\tilde{\gamma}(x)|\lesssim|x|^{-\left(\alpha-\zeta/2\right)}=|x|^{-\eta/2}$. Hence, $\tilde{\gamma}$ is a $(\eta/2)$-amplitude spectral density. Now, by definition of $\tilde{\gamma}$,
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}$ $=\|\tilde{\gamma}(\mathcal{L})\mathcal{L}^{-\zeta/2}\mathcal{W}-\tilde{\gamma}(\mathcal{L}_{h})\mathcal{L}_{h}^{-\zeta/2}P_{h}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
so that the triangle inequality yields $\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq E_{1}+E_{2}$ where we take
$E_{1}$ $=\|\big{(}\tilde{\gamma}(\mathcal{L})-\tilde{\gamma}(\mathcal{L}_{h})P_{h}\big{)}\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))},$
$E_{2}$ $=\|\tilde{\gamma}(\mathcal{L}_{h})\big{(}P_{h}\mathcal{L}^{-\zeta/2}\mathcal{W}-\mathcal{L}_{h}^{-\zeta/2}P_{h}\mathcal{W}\big{)}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}.$
For the term $E_{1}$, we first note that $\zeta=d/2+2(\alpha-d/2-\varepsilon_{h})>d/2$ since $\varepsilon_{h}<\alpha-d/2$. Therefore, we can use [6, Lemma 4.2] and the fact that $h$ (and therefore $\varepsilon_{h}$) is upper-bounded to deduce that
$\|\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}\leq\frac{\zeta}{\zeta-d/2}\lesssim 1.$
In particular, $\mathcal{L}^{-\zeta/2}\mathcal{W}$ is in $L^{2}(\Omega;L^{2}(\mathcal{M}))$. Let then $p=\min\{\alpha-d/2;1\}-2\varepsilon_{h}$. In particular, we have $p\in(0,1)$ since $p\leq 1-2\varepsilon_{h}<1$ and $p>\min\{\alpha-d/2;1\}-\min\{\alpha-d/2;d/2\}\geq 0$. Moreover, we have $\zeta/2-p=d/4+(\alpha-d/2)-\min\{\alpha-d/2;1\}+\varepsilon_{h}\geq d/4+\varepsilon_{h}>d/4$. Hence, using once again [6, Lemma 4.2], we can conclude that
$\|\mathcal{L}^{p}\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$ $=\|\mathcal{L}^{-(\zeta/2-p)}\mathcal{W}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
$\lesssim\frac{\zeta/2-p}{\zeta/2-p-d/4}\lesssim\begin{cases}|\log h|,&\text{if }\alpha-d/2\leq 1\\ 1,&\text{if }\alpha-d/2>1\end{cases}$
where we used the definition of $\varepsilon_{h}$ to derive the last inequality. Hence, we can apply Proposition 2.4 to get
$E_{1}^{2}$ $=\operatorname{\amsmathbb{E}}[\|\big{(}\tilde{\gamma}(\mathcal{L})-\tilde{\gamma}(\mathcal{L}_{h})P_{h}\big{)}\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\mathcal{M})}^{2}]$
$\lesssim C_{\eta/2+p}(h)^{2}h^{4\min\{\eta/2+p;1\}}\operatorname{\amsmathbb{E}}[\|\mathcal{L}^{p}\mathcal{L}^{-\zeta/2}\mathcal{W}\|_{L^{2}(\mathcal{M})}^{2}]$
$\lesssim C_{\eta/2+p}(h)^{2}h^{4\min\{\eta/2+p;1\}}K^{\prime}_{\alpha}(h)^{2},$
where we take on the one hand, $C_{\eta/2+p}(h)=|\log h|$ if $\eta/2+p\leq 1$, and $C_{\eta/2+p}(h)=1$ otherwise, and on the other hand $K^{\prime}_{\alpha}(h)=|\log h|^{1/2}$ if $\alpha\leq 1+d/2$ and $K^{\prime}_{\alpha}(h)=1$ otherwise.
Finally, note that according to Lemma A.1, $\eta/2+p=\min\{\alpha-d/2;1\}+d/4-\varepsilon_{h}\geq\min\{\alpha-d/4;1\}-\varepsilon_{h}$ which implies that $h^{4\min\{\eta/2+p;1\}}\leq h^{4(\min\{\alpha-d/4;1\}-\varepsilon_{h})}=h^{4\min\{\alpha-d/4;1\}}\exp(2\min\{d/2;\alpha-d/2\})$. Besides, we note that if $\alpha>1+d/2$, then $\eta/2+p=1+d/4-\varepsilon_{h}>1$ (since by construction $\varepsilon_{h}<d/4$), and therefore $C_{\eta/2+p}(h)=1$. We can therefore conclude that
$E_{1}\lesssim K_{1}(h)h^{2\min\{\alpha-d/4;1\}}$
where $K_{1}(h)=|\log h|^{3/2}$ if $1<\alpha\leq 1+d/4+\varepsilon_{h}$, and $K_{1}(h)=|\log h|^{1/2}$ if $21+d/4+\varepsilon_{h}<\alpha\leq 1+d/$ 2 and $K_{1}(h)=1$ if $\alpha>1+d/2$.
For $E_{2}$, we first use the definition of the projection operator $P_{h}$ and the self-adjointness of $\mathcal{L}$ and $\mathcal{L}_{h}$ (which is a consequence of the definition of the associated bilinear form $\mathsf{A}_{\mathcal{M}}$)
$E_{2}^{2}$ $=\bigg{\|}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})\bigg{(}\big{(}P_{h}\mathcal{L}^{-\zeta/2}\mathcal{W},e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}-\big{(}\mathcal{L}_{h}^{-\zeta/2}P_{h}\mathcal{W},e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}\bigg{)}e_{i}^{h}\bigg{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
$=\bigg{\|}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})\bigg{(}\big{(}\mathcal{L}^{-\zeta/2}\mathcal{W},e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}-\big{(}P_{h}\mathcal{W},\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}\bigg{)}e_{i}^{h}\bigg{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
$=\bigg{\|}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})\bigg{(}\big{(}\mathcal{W},\mathcal{L}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}-\big{(}\mathcal{W},\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}\bigg{)}e_{i}^{h}\bigg{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}$
$=\bigg{\|}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})\big{(}\mathcal{W},\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}e_{i}^{h}\bigg{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}))}^{2}.$
Using the orthonormality of the basis $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ and the definition of the white noise $\mathcal{W}$, we conclude that
@@
$E_{2}^{2}$ $=\operatorname{\amsmathbb{E}}\bigg{[}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}|\big{(}\mathcal{W},\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\big{)}_{L^{2}(\mathcal{M})}|^{2}\bigg{]}$
$=\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\|\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}^{2}.$
On the one hand, using Proposition 2.4 with $f=e_{i}^{h}\in S_{h}^{\ell}$ ($1\leq i\leq N_{h}$), which has norm $\|e_{i}^{h}\|_{L^{2}(\mathcal{M})}=1$, we have
$\|\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}\lesssim C_{\zeta}(h)h^{2\min\{\zeta/2;1\}},\quad\text{where}\quad C_{\zeta}(h)=\begin{cases}1,&\zeta>2,\\ |\log h|,&\zeta\leq 2.\end{cases}$
Note that by definition of $\zeta$ and using Lemma A.1, we have $\min\{\zeta/2;1\}\geq\min\{\alpha-d/4;1\}-\varepsilon_{h}>0$ (since $\varepsilon_{h}\in(0,\alpha-d/4)$), which in turn gives $h^{2\min\{\zeta/2;1\}}\leq h^{2(\min\{\alpha-d/4;1\}-\varepsilon_{h})}\lesssim h^{2\min\{\alpha-d/4;1\}}$. Hence, we can conclude that
$\|\mathcal{L}^{-\zeta/2}e_{i}^{h}-\mathcal{L}_{h}^{-\zeta/2}e_{i}^{h}\|_{L^{2}(\mathcal{M})}\lesssim C_{\zeta}(h)h^{2\min\{\alpha-d/4;1\}}$
and therefore, $E_{2}^{2}\lesssim C_{\zeta}(h)^{2}h^{4\min\{\alpha-d/4;1\}}\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}$. On the other hand, the term $\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}$can be bounded using Equation 12 and Equation 6 which yield that
$\sum_{i=1}^{N_{h}}\tilde{\gamma}(\lambda_{i}^{h})^{2}\lesssim\sum_{i=1}^{N_{h}}(\lambda_{i}^{h})^{-\eta}\lesssim\sum_{i=1}^{N_{h}}i^{-2\eta/d}\lesssim\int_{1}^{\infty}t^{-2\eta/d}=\frac{1}{2\eta/d-1}\lesssim|\log h|,$
since $12\eta/d=1+4\varepsilon_{h}/d>$ 1 and by definition of $\varepsilon_{h}$. Hence, we can conclude that
$E_{2}$ $\lesssim C_{\zeta}(h)|\log h|^{1/2}h^{2\min\{\alpha-d/4;1\}}\lesssim K_{2}(h)h^{2\min\{\alpha-d/4;1\}},$
where $K_{2}(h)=C_{\zeta}(h)|\log h|^{1/2}$, i.e., $K_{2}(h)=|\log h|^{3/2}$ if $1<\alpha\leq 1+d/4+\varepsilon_{h}$, and $K_{2}(h)=|\log h|^{1/2}$ if $\alpha>1+d/4+\varepsilon_{h}$.
Hence, by putting together the estimates for $E_{1}$ and $E_{2}$, and noting that $K_{1}(h)\leq K_{2}(h)$, we can conclude that if $\alpha>1$, then
$\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\leq E_{1}+E_{2}\lesssim K_{2}(h)h^{2\min\{\alpha-d/4;1\}}.$
In particular note that for any $\alpha>1$, we can clearly upper-bound $K_{2}(h)$ by $K_{2}(h)\lesssim|\log h|^{3/2}$. But if moreover $\alpha\geq 1+d/2$, we have y definition of $\varepsilon_{h}$, $1+d/4+\varepsilon_{h}<1+d/4+\min\{d/2;\alpha-d/2\}/2\leq 1+d/2\leq\alpha$, and therefore $K_{2}(h)=|\log h|^{1/2}$. Hence, we can conclude that if $\alpha>1$, $\|\mathcal{Z}-\tilde{\mathcal{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}))}\lesssim K^{\prime}_{2}(h)h^{2\min\{\alpha-d/4;1\}}$ where $K^{\prime}_{2}(h)=|\log h|^{3/2}$ if $21<\alpha<1+d/$ 2 and $K^{\prime}_{2}(h)=|\log h|^{1/2}$ if $\alpha\geq 1+d/2$. And finally gathering the estimates obtained in the two cases $\alpha\leq 1$ and $\alpha>1$, we obtain the result stated in the lemma. ∎
@@
For the error between $\big{(}\tilde{\mathcal{Z}}_{h}\big{)}^{-\ell}$ and $\widehat{\mathsf{Z}}_{h}$ we get the following estimate, inspired by [6, Lemma 4.4].
@@
Let
$\mathcal{E}_{h}=\big{\|}\left(\tilde{\mathcal{Z}}_{h}\right)^{-\ell}-\widehat{\mathsf{Z}}_{h}\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}=\big{\|}\left(\gamma(\mathcal{L}_{h})\widetilde{\mathcal{W}}\right)^{-\ell}-\gamma(\mathsf{L}_{h})\mathsf{P}_{h}(\sigma\widetilde{\mathcal{W}}^{-\ell})\big{\|}_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}.$
We note that $\widetilde{\mathcal{W}}$ is an $S_{h}^{\ell}$-valued random variable. Therefore, we can apply Equation 26 to realizations of $\widetilde{\mathcal{W}}$, and take the expectation on both sides to get
@@
We then distinguish two cases. First, if $d=1$, then we note that $\min\{\alpha+d/4;1\}\geq\alpha+d/4>d/2$ since $\alpha>d/4$. Besides, [6, Lemma 4.2] yields that for all $r\in(d/2,2)$, there is a constant $C>0$ such that
@@
If now $d=2$, since $\alpha+d/4>d/2=1$, we have $\min\{\alpha+d/4;1\}=1$. We then note that by the proof of [6, Lemma 4.2], for any $\varepsilon>0$
@@
Let $h\in(0,h_{0})$. It holds that there is a constant such that
$\|\mathsf{Z}_{h}-\widehat{\mathsf{Z}}_{h}\|_{L^{2}(\Omega;L^{2}(\mathcal{M}_{h}))}\lesssim h^{2}.$
@@
where [11, Lemma 4.1] was applied in the final step. We conclude that